OpenStudy (anonymous):

What is the y-intercept of the line 3x + 4y = 16? A. B. 16 C. 4 D. 3

OpenStudy (anonymous):

3x+4y=16 4y=-3x+16 y=-3/4 + 4 hence y=mx+b and b is the y-intercept, so 4 is your b

OpenStudy (anonymous):